An angle's measure is 6 degrees more than 3 times the meausre of its compliment. Find the measure of the angle.
Answer by cleomenius(959)

x + 6 + 3x = 90
4x = 84
x = 21 degrees, the smaller angle.
3(21) + 6 = 69 degrees.
This is the complement.
69 + 21 = 90 degrees.
cleomenius.